# Charged Scalar Field at Future Null Infinity via Nonlinear Hyperboloidal Evolution

### v2

@@ Abstract
Quasinormal modes and power-law late-time decay tails of a charged scalar field in a charged black hole background have been studied, but never in the fully non-linear regime, as far as we know. In this paper, we study the dependence of these properties on the charges of scalar field and black hole. For the quasinormal modes, a fit of the spherical fundamental mode is shown for the purely uncharged case and compared to the charged one. We also see for the first time the transition from gravitational decay to pure electromagnetic decay, and show disagreement with the oscillation frequency between real and imaginary parts of the scalar field prescribed in the literature. Full non-linear evolutions of hyperboloidal slices in spherical symmetry were used to obtain our results, allowing for the extraction of signals at future null infinity.
@@ I Introduction
The seminal work price_nonspherical_1972 motivated a still-growing interest in QNMs and power-law decay tails. For the case of a real scalar field, the power-law decay comes from the backscattering of outgoing waves of the scalar field at late-times, due to the effective potential created by the presence of a black hole. Being $l$ the mode of the scalar field $\phi$, in the limit of timelike infinity the expectation is that
$\phi\propto t^{-2l-2}\,.$ (1)
Shortly afterwards, Ref. bicak_gravitational_1972 extended the late-time power-laws to the charged scalar field collapsing into a RN black hole. The same power-law as (1) was predicted, but with the signal dominated by slower decaying primary waves.
About two decades later, Ref. gundlach_late-time_1994 studied massless scalar field perturbations in fixed Schwarzschild and RN spacetimes, but this time also focusing on future null infinity – the collection of the end-points of future-directed null geodesics. Not long after, Hod and Piran  [^1] We mention the authors’ names for ease of comparison with their results in Figs. 7 and 8. published a three-paper investigation on the late-time evolution of charged gravitational collapse and decay of charged scalar hair hod_late-time_1998; hod_late-time_1998_II; hod_late-time_1998_III. In Ref. hod_late-time_1998 they agreed with Ref. bicak_gravitational_1972, but only for the regime of $qQ\ll 1$, where $q$ is the scalar field’s charge parameter and $Q$ is the black hole’s charge. The second paper then corrects the expected power-law for general values of $qQ$, reading
$\phi\propto t^{-2\beta+2}\,$ (2)
at timelike infinity. The quantity $\beta$ is defined as,
@@
At future null infinity, they also predict that,
$\phi\propto u^{-(\beta+1-iqQ)}\,$, (4)
where $u$ denotes the retarded time. The difference between (1) and (3) is that as the parameter $qQ$ increases, the electromagnetic coupling between the charged scalar field’s real and imaginary parts becomes stronger and the QNMs that we expect to see start to be dominated by electromagnetism. The same goes for the tail decay.
@@ V Late-time behavior
To check the late-time behavior of the scalar field’s perturbation of a RN black hole (after the QNM ringing), we now use the initial data for $\bar{c}_{\Pi}$ in (27). The scalar field’s charge is set to 1.0. Regarding the charge of the black hole $Q$, we have ran several simulations, each with a different value of $Q$, to see the effect of the parameter $qQ$ bicak_gravitational_1972; konoplya_massive_2013; hod_late-time_1998 on the exponent of the power-law regarding the decay. The results are shown in Fig. 5. The relevant quantity to look at is the amplitude of the scalar field (20). For low values of $qQ$ (i.e., the top left graph), both the real and imaginary parts enter a power-law decay, mostly dominated by the real part. However, as we increase $qQ$, the real and imaginary parts continue to oscillate one around the other, presumably due to the increased electric coupling between each other. Even during their oscillation, their amplitude decreases with a fixed power-law tail. Note also that the frequency of oscillation during the tail regime increases with $qQ$.
@@ VI Tail Regime
$p_{\mathrsfs{I}^{+}}=\text{Re}(-\beta)-1,\quad p_{i^{+}}=\text{Re}(-2\beta)+2\,$, (32)
@@
For $qQ=0.6$ and $qQ=1.0$ both signals oscillate around p=0.5p=-0.5 and p=1.0p=-1.0 for $\mathrsfs{I}^{+}$ and $r\approx 0.75$, respectively, meaning that the maximum real part of $\beta$ has been reached (check (32) and (3)). To understand better how the exponent changes, we considered more values of $qQ$. To get precise values of $p$, we did a linear regression on $d\log|\phi|/d\log t$ between $t=100$ and $t=150$ and the slope of each regression corresponds to the $p$ of that same simulation. The results are plotted in Fig. 7, where they are compared to previous analytical results in the literature.
@@
Figure 7: Comparison of the results from our simulation (the triangles) with the analytical results coming from Konoplya and Zhidenko konoplya_massive_2013 and from Hod and Piran hod_late-time_1998. As was mentioned in the main body of the text, the red lines correspond to the expected decay of the scalar field at $i^{+}$. The red points, however, correspond to the decay of the scalar field at a given radius $r\approx 0.75$ and at finite time, which is not $i^{+}$. Nonetheless, as the simulation time increases, the decay we expect to see at $r<1$ should approach asymptotically the one at $i^{+}$.
The most relevant conclusion here is that expressions (4) and (3) describe well how $p$ changes with $qQ$, except for $qQ\ll 1$, where the better approximation is the same as in the uncharged case. There is a transition phase between both regimes, which indicates the change from gravitational to electromagnetic effects. As far as we know, this is the first time a continued comparison has been shown of both regimes of the scalar field’s tail decay. Although this effect has been addressed in the literature, there has not been a clear understanding of how this transition happens. It would be interesting to explore better how to construct a tail decay that is partially affected by the gravitational curvature and by the electromagnetic coupling.
We also wanted to see the effect of $qQ$ in the oscillation frequency of the real and imaginary parts of the scalar field at $\mathrsfs{I}^{+}$. For $qQ<1/2$, (4) implies that the frequency should increase linearly with this charge parameter. For $qQ>1/2$, we have to take into account the imaginary part that comes from $\beta$. To compare the analytical results with our simulation, we applied a Fourier transform to the real part (or equivalently to the imaginary part) of the scalar field during the tail regime. When plotting our results against the prediction (4), there seemed to be an order of magnitude of difference, namely, that for $qQ<1/2$, the frequencies coming from our simualation were exactly 10 times smaller. Therefore, in Fig. 8, we show the imaginary part of (4) divided by 10, the frequencies of the real part of the scalar field as $qQ$ increases, and a blue line that corresponds to Frequency =qQ/10=qQ/10. This blue line is an empirical guess that suits better the data than the prediction.Thus, while for $p$ the theoretical predictions were in agreement with our simulations, for the frequency analysis we see a clear divergence especially for $qQ>1/2$.
@@
Figure 8: Frequency comparison of the results from our simulation (the triangles) with one tenth of the analytical results coming from Hod and Piran hod_late-time_1998, (4), (the black line). The blue line corresponds to the function $f=Qq/10$, an empirical guess that shows better agreement.
@@ VII Conclusions
With this work, we took advantage of hyperboloidal slices to gain access to $\mathrsfs{I}^{+}$ efficiently to understand the behavior of a charged scalar field perturbation of a RN black hole. This allowed us to study the QNM ringing of the scalar field in Schwarzschild and RN spacetimes and found it only possible to fit the spherical fundamental mode to the former case. The latter showed but one oscillation, making it difficult to construct a reliable fit. We propose that the reason for the almost complete absense of oscillations is due to the self interaction of the charged scalar field, which picks up as soon as the imaginary part is excited by the presence of the charged black hole. Regarding the tail regime, although our results agree very well with Ref. hod_late-time_1998_III, there is a transition region where the decay is affected both by the black hole’s curvature and also by the electromagnetic coupling that to this day seems to have been unadressed, to the best of our knowledge. The fact that this transition occurs around $Qq\sim 10^{-3}$, means that, contrary to the point-of-view that electrically charged black holes might be astrophysically irrevelant shapiro_black_1983, even if a black hole has a very small amount of charge, it may have a noticeable impact on the late-time behavior of perturbations.
We also analysed the frequency of oscillation of the real and imaginary parts during the tail regime and found a linear correlation between $Qq$ and the frequency with which they oscillate. For $Qq>1/2$, the expected frequency seems to deviate from the analytical predictions in Ref. hod_late-time_1998_III.
@@ Acknowledgements
The authors would like to thank Christian Peterson Bórquez for useful advice regarding the ringing of QNM in hyperboloidal slices and to Adrien Kuntz and Nicola Franchini for the help finding important references for this work. The acknowledgements extend naturally to Vitor Cardoso and again to Adrien for very useful comments on the manuscript. We also thank the PRD referee for spotting a mistake in the equations that had gone undetected until now.

### v4

@@ Abstract
Quasinormal modes and power-law late-time decay tails of a charged scalar field on a charged black hole have been considered in the literature, and here we use the hyperboloidal formalism to present an independent comparison to prior non-linear results. For the quasinormal modes we show a fit of the spherical fundamental mode for the purely uncharged case and compare it to the charged one. We obtain good agreement with prior prescriptions for the oscillation frequency between real and imaginary parts of the scalar field, as well as for the exponent of the power-law decay. Full non-linear evolutions of hyperboloidal slices in spherical symmetry were used to obtain our results, allowing for the extraction of signals at future null infinity. [^1] Note: this version of the paper has been revised to take into account the comments in Hod:2025miv.
@@ I Introduction
The seminal work price_nonspherical_1972 motivated a still-growing interest in QNMs and power-law decay tails. For the case of a real scalar field, the power-law decay comes from the backscattering of outgoing waves of the scalar field at late-times, due to the effective potential created by the presence of a black hole. Being $l$ the mode of the scalar field $\phi$, in the limit of timelike infinity the expectation is that price_nonspherical_1972
$\phi\propto t^{-{(2l+3)}}\,.$ (1)
Shortly afterwards, Ref. bicak_gravitational_1972 analyzed the late-time power-laws of a neutral scalar field on an electrically charged star collapsing into a RN black hole. The same power-law as (1) was predicted, but with the signal dominated by slower decaying primary waves.
About two decades later, Ref. gundlach_late-time_1994 studied massless scalar field perturbations in fixed Schwarzschild and RN spacetimes, but this time also focusing on future null infinity – the collection of the end-points of future-directed null geodesics. Not long after, Hod and Piran  [^2] We mention the authors’ names for ease of comparison with their results in Figs. 7 and 8. published a three-paper investigation on the late-time evolution of charged gravitational collapse and decay of charged scalar hair hod_late-time_1998; hod_late-time_1998_II; hod_late-time_1998_III. In Ref. hod_late-time_1998 they agreed with Ref. bicak_gravitational_1972, but only for the regime of $qQ\ll 1$, where $q$ is the scalar field’s charge parameter and $Q$ is the black hole’s charge. The second paper then corrects the oscillatory power-law tail for general values of $qQ$, reading
$\phi\propto{e^{i\frac{qQ}{r_{+}}t}}\ t^{-{(}2\beta+2{)}}\,$ (2)
at timelike infinity, with $r_{+}=M+\sqrt{M^{2}-Q^{2}}$ the event horizon of the RN black hole. The quantity $\beta$ is defined as
@@
At future null infinity, they also predict that
$\phi\propto{e^{i\frac{qQ}{r_{+}}t}}u^{-(\beta+1-iqQ)}\,$, (4)
where $u$ denotes the retarded time , and is related to $t$ as $t=(u+v)/2$ with $v$ the advanced time. The difference between (1) and (2) is that as the parameter $qQ$ increases, the electromagnetic coupling between the charged scalar field’s real and imaginary parts becomes stronger and the QNMs that we expect to see start to be dominated by electromagnetism. The same goes for the tail decay. Note that (2) does not reduce to (1) for $qQ\to 0$, because this limit does not commute with $t\to\infty$, as understood in hod_late-time_1998. Finally, the numerical non-linear results presented in hod_late-time_1998_III confirm predictions from the earlier papers.
@@ V Late-time behavior
To check the late-time behavior of the scalar field’s perturbation of a RN black hole (after the QNM ringing), we now use the initial data for $\bar{c}_{\Pi}$ in (27). The scalar field’s charge is set to 1.0. Regarding the charge of the black hole $Q$, we have ran several simulations, each with a different value of $Q$, to see the effect of the parameter $qQ$ bicak_gravitational_1972; konoplya_massive_2013; hod_late-time_1998 on the exponent of the power-law regarding the decay. The results are shown in Fig. 5. The relevant quantity to look at is the amplitude of the scalar field (20). For low values of $qQ$ (i.e., the top left graph), both the real and imaginary parts enter a power-law decay, mostly dominated by the real part. However, as we increase $qQ$, the real and imaginary parts continue to oscillate one around the other, presumably due to the increased electric coupling between each other. Even during their oscillation, their amplitude decreases with a fixed power-law tail. Note also that the frequency of oscillation during the tail regime increases with $qQ$, see Fig. 8 and corresponding explanation for further details.
@@ VI Tail Regime
$p_{\mathrsfs{I}^{+}}=\text{Re}(-\beta)-1,\quad p_{i^{+}}=\text{Re}(-2\beta){-}2\,$, (32)
@@
For $qQ=0.6$ and $qQ=1.0$ both signals oscillate around p=0.5p=-0.5 and p=1.0p=-1.0 for $\mathrsfs{I}^{+}$ and $r\approx 0.75$, respectively, meaning that the maximum real part of $\beta$ has been reached (check (32) and (3)). To understand better how the exponent changes, we considered more values of $qQ$. To get precise values of $p$, we did a linear regression on $d\log|\phi|/d\log t$, whose slope provides the value of $p$. For $qQ\gtrsim 0.1$ performing the regression between $t=200$ and $t=500$ is valid, as the asymptotic regime is already reached at those times. However, in the limit $qQ\ll 1$ the electromagnetic contribution dominates asymptotically for $t\gg\frac{M}{qQ}$ hod_late-time_1998; hod_late-time_1998_II, so that simulations need to be evolved for longer to reach the desired regime. In Fig. 7 we show the results of several simulations, where we have fixed $M=q=1$, and we change the value of $Q$. This makes the overall factor $qQ$ change exactly in the same way as $Q$. The asymptotic exponents (i.e., respecting the condition $t\gg\frac{M}{qQ}$) extracted from our simulation are plotted in triangles. Solid lines represent the analytical predictions as shown in Ref. hod_late-time_1998_III, while dashed lines correspond to the predictions shown in Ref. konoplya_massive_2013. We also show the values of $p$ for the case when $Q=0.001$ estimated between $t=200$ and $t=500$, represented through two crosses in the same color scheme. For this small value of $qQ$, these times correspond to the intermediate regime instead of the asymptotic one – to obtain $p$ for the latter we ran the simulation until $t=6000$. It is interesting to note that initially, when the charged scalar field in the presence of a charged black hole enters the tail regime, it initially behaves as the neutral scalar field. This happens because the amplitude of the charged scalar field is mostly dominated initially by the real or complex part of the field, depending whether we perturb the former or the latter in the initial data. Only after some time, when both parts become comparable, does the asymptotic exponent corresponding to the charged scalar field appear. The main conclusion from this comparison is that our simulations coincide well with previous analytical results in the literature. Furthermore, in Fig. 7, we show not only the values of $p$ at $\mathrsfs{I}^{+}$ (in blue), but also the values along a timelike hypersurface corresponding to the compactified value of $r=0.751$ (in red). The main idea behind comparing the values of $p$ along these hypersurfaces to the values of $p$ as expected at $i_{+}$ is that as time progresses onwards, the observer moves ever closer to the position of $i_{+}$. It will never reach it, but it asymptotes to it. This is exactly what we see in the red quantities in Fig. 7.
@@
Figure 7: Comparison of the results from our simulation (the triangles and crosses) with the analytical results coming from Konoplya and Zhidenko konoplya_massive_2013 and from Hod and Piran hod_late-time_1998. Blue lines and symbols correspond to values at $\mathrsfs{I}^{+}$. As mentioned in the main body of the text, the red lines correspond to the expected decay of the scalar field at $i^{+}$. The red triangles, however, correspond to the decay of the scalar field at a given radius $r\approx 0.75$ and at finite time, which is not $i^{+}$. Nonetheless, as the simulation time increases, the decay we expect to see at $r<1$ should approach asymptotically the one at $i^{+}$.
We also wanted to see the effect of $qQ$ in the oscillation frequency of the real and imaginary parts of the scalar field at $\mathrsfs{I}^{+}$. The oscillatory part of (4) is
$\phi_{\textrm{oscil}}=e^{i\varphi_{\textrm{oscil}}}$ $=e^{i\frac{qQ}{r_{+}}t}u^{i\left[qQ-\text{Im}(\beta)\right]}$
$=e^{i\frac{qQ}{2\,r_{+}}u}e^{i\left[qQ-\text{Im}(\beta)\right]\log(u)}$
$=e^{i\left\{\frac{qQ}{2\,r_{+}}u+\left[qQ-\text{Im}(\beta)\right]\log(u)\right\}},$
so that its $u$-dependent angular frequency is
$\omega=\frac{d\varphi_{\textrm{oscil}}}{du}=\frac{qQ}{2\,r_{+}}+\left[(qQ-\text{Im}(\beta)\right]\frac{1}{u}.$ (33)
To compare the analytical results with our simulation, we applied a Fourier transform to the real part (or equivalently to the imaginary part) of the scalar field during the tail regime. This analysis is shown by the red triangles in Fig. 8. The predicted frequency (33) (divided by $2\pi$ for ease of comparison) is evaluated at $u=500$, i.e., the same value of the retarded time as that at which we are extracting the frequency. This is because our hyperboloidal time $t$ at $\mathrsfs{I}^{+}$ and $u$ are comparable in the asymptotic regime. As $u$ increases, the shape of (33) becomes ever more dominated by the factor $qQ/(2r_{+})$. We can therefore see that there is a good match between our results and the prediction by (33), except for large values of $qQ$ close to extremality, where the divergence becomes quite stark and the data points from our simulation deviate. Understanding why this happens is left for future work.
@@
Figure 8: Frequency comparison of the results from our simulation (the triangles) with the analytical results coming from Hod and Piran hod_late-time_1998, (4), where the black line corresponds to the frequency given by (33) divided by $2\pi$ and evaluated at $u=500$ (visually indistinguishable from the same expression evaluated at $u=2000$).
@@ VII Conclusions
With this work, we took advantage of hyperboloidal slices to gain access to $\mathrsfs{I}^{+}$ efficiently to understand the behavior of a charged scalar field perturbation of a RN black hole. This allowed us to study the QNM ringing of the scalar field in Schwarzschild and RN spacetimes and found it only possible to fit the spherical fundamental mode to the former case. The latter showed but one oscillation, making it difficult to construct a reliable fit. We propose that the reason for the almost complete absense of oscillations is due to the self interaction of the charged scalar field, which picks up as soon as the imaginary part is excited by the presence of the charged black hole. Regarding the tail regime, our results agree very well with Ref. hod_late-time_1998_III. Contrary to the point-of-view that electrically charged black holes might be astrophysically irrevelant shapiro_black_1983, even if a black hole has a very small amount of charge, it may have a noticeable impact on the late-time behavior of perturbations.
We also analysed the frequency of oscillation of the real and imaginary parts during the tail regime and found good agreement with the analytical predictions in Ref. hod_late-time_1998_III, although differences appear towards the extremal case.
@@ Acknowledgements
The authors would like to thank Christian Peterson Bórquez for useful advice regarding the ringing of QNM in hyperboloidal slices and to Adrien Kuntz and Nicola Franchini for the help finding important references for this work. The acknowledgements extend naturally to Vitor Cardoso and again to Adrien for very useful comments on the manuscript. We also thank the PRD referee for spotting a mistake in the equations that had gone undetected until now. Finally, we are indebted to Shahar Hod for pointing out several inaccuracies in an earlier version of our paper.